3(a). You want to retire 50 years from now, and you want access to ten million dollars ($10,000,000) when you do. You decide to put some money towards buying a certificate of deposit that’ll be worth 10 mil in 50 years. The CD has an annual interest rate of 8%, and it compounds quarterly. How much do you have to put into this deposit?3

(b). Let’s say you put an equal amount of money into a simple interest account with 16% annual interest. Is this a better investment? Determine this by calculating how much money will be in this account after 50 years
Answer #1

b)Quarterly interest rate= 8%/4=2%

Total number of period(quarter) in 50 year=50*4=200

Say, x is the amount of money that is required

So, X*(1+2%)^200=10000000 by solving X=1,90,531

So, the amount of money that is required to deposit =$1,90,531

b) If you put the money in simple interest account, then per year interest=190531*16%=$30484.96

So, in 50 year total money will become= 30484.96*50+1,90,531=$1,714,779

As in this case the money is less the first option is better.